< ⓘ Assignrment Score 1167/1700 Resources Give Up? Feedback Question 9 of 17 > Attempt 1 ▼ What are the magnitude Au and angle 0 of the change in velocity if the initial velocity is 65.0 ms south and the final velocity is 80.0 m/s west? Let the angle θ be measured relative to the east direction (eE-:: 90° meahs Δυ points north, θ-180. means Aö points west, etc.). Δν 103.078 m/s
Answer #1

answer) here initial velocity is

vo=-65j ( pointing south which is negative y direction)

vf=-80i( pointing west which is negative x direction)

so \Deltav=vf-vo=-80i+65j

magnitude=\sqrt{}(-802)+652=103 m/s

so the answer is 103 m/s or 103.1m/s

and angle =tan-1(65/-80)=-39o

so angle=-39o or -39.1o.( if the answer comes out wrong try 39o or 39.1o and omit the negative sign)